OpenStudy (anonymous):

Find two consecutive integers whose sum is 113. Which of the following equations could be used to solve the problem?

OpenStudy (anonymous):

2x = 113 2x + 1 = 113 2x + 2 = 113 x^2 + 1= 113

OpenStudy (anonymous):

2x + 1 = 113

OpenStudy (anonymous):

You have two integers, you can represent the first one as x, since they are consecutive the next one is x+1. Add them together you get 2x+1, which = 113.

OpenStudy (anonymous):

okoy doky

OpenStudy (apoorvk):

Well, two consecutive integers right? so let one integer be 'x'. Since they are consecutive integers, the other integer would be 'x+1' now their sum is '113'. So, x + x+1 = 113 Can you solve this^?

OpenStudy (anonymous):

56 and 57

OpenStudy (anonymous):

2x+1=113

Parth (parthkohli):

Think of this like this. If there are two consecutive numbers(integers), then we'll have a number and a number one more than that. \(x, x + 1\) can be said as the numbers needed. We have that the sum is 113. \( \color{Black}{\Rightarrow x + (x + 1) = 113 }\) Which can be further simplified to: \(2x + 1 = 113\) by combining the x's.

OpenStudy (apoorvk):

rather, you don't need to solve this equation here - just add the 2 'x's you get 2x + 1 = 113.

OpenStudy (anonymous):

yes i like thisbetter